#1
DATE: 5/5/93
PRIORITY: 2
contact Dave Carter on fibre option with ISIS. what is status ?
STATUS: 8/5/93: got draft of manual/description of fibres.
had a quick look at the system during setup.
10/5/93: Dave Carter explained the system to me. See notes.
RESULT: released
#2
DATE: 5/5/93
PRIORITY: 2
contact Charles Jenkins on (i) cross disperser situation, and
(ii) anamorphotic lens
STATUS: 10/5/93: Dave Carter says one cross disperser will be delivered
this summer (for the blue arm). For the red arm it will depend
on interest expressed by the community whether it will be ordered.
The anamorphotic lens will probably never be ordered.
RESULT: released
#3
DATE: 5/5/93
PRIORITY: 2
make inventory of all ISIS stuff, clean out cupbord in instrument
store and label all items
STATUS: 16/11/93: looks okay
RESULT: dismissed
#4
DATE: 5/5/93
PRIORITY: 1
make quality control document
STATUS: 26/5/93: done
RESULT: released
#5
DATE: 5/5/93
PRIORITY: 2
polarisation:
(i) check out FOS polarisation
(ii) observe polarisation standards
(iii) get circular polarisation going
STATUS: 26/7/93: (i) done, seems to work okay
(ii) ongoing; obtained several good standards now but need more
to monitor and get more pol. standards
(iii) is working okay !
RESULT: released
#6
DATE: 8/5/93
PRIORITY: 2
The ISIS manual needs to be updated with new commands for the autoguider
control and the fibres option. Dave Carter will consult LPO for input.
STATUS: will be in press soon
RESULT: in hand
#7
DATE: 13/5/93
PRIORITY: 2
dichroics efficiency test (test all dichroics).
Red arm: 158 grating, cenwave = dichr cutoff + 600 A
tungsten lamp exposure with and without dichroic in beam + arc expo.
Blue arm: 158 grating, cenwave = dichr cutoff - 600 A
absolutely dark conditions (long exposures !).
Deuterium lamp exposure with dichroic and with mirror in beam + arc expo.
STATUS: 26/8/93: got data. reduction underway (Uli)
18/12/93: 2 red dichroics done. Blue ones need better exposures, but
since we're getting new ones we'll put this on hold.
RESULT: released
#8
DATE: 13/5/93
PRIORITY: 2
ask for more dichroic slides and mirrors mounted on them
from enhancements budget ?
STATUS: 6/93: a set of slides is being made and new dichroics are on order
3/94: new dichroics delivered and mounted.
RESULT: released.
#9
DATE: 15/5/93
PRIORITY: 2
Janet Sinclair has had some requests to extend the ISIS arc map
set. After some discussion we propose to add:
- 150 grating maps (Cune and CuAr) with both arms
- 1200 grating maps ( " " ) red arm only.
in that order of priority. Could you help to get these data?
Suggested setups:
Slit 2 pixels. 200 pixels window (?) to improve S/N
150 grat, blue - wave_range 2900 - 5840A; centre thus 4374.
150 grat, red - wave_range_1 5000-8350, centre thus 6675
2 8000-11000, 9675.
1200grat, red only. We get about 470A per spectrum. Suggest
start from 6000A and work up to 11,000A, with a decent overlap
between spectra (eg, 15%)...Janet is keen on this.
This is over 20 spectra (2 lamps)...quite a lot.
By the way, Janet suggests not using an order-sorting filter; she
finds it useful to identify the 2nd order lines.
Dave mentioned that he might be able to get these data; otherwise
can a LP SA help in daytime? Janet will reduce the data.
STATUS: 10/7/94 done
RESULT: released
#10
DATE: 15/5/93
PRIORITY: 2
image of slit on TV has slipped down. Investigate.
STATUS: Sue W. will fix this during Cass rot standdown in June.
2/6/93: alligned slit view mirror and scale 12 lens by looking with
TV at fake slit ('top-hat' with mirror under 7.5 degrees from vertical,
as slit, which can be attached to A&G box).
Also cleaned optics with air spay and peel-off plasic spray.
RESULT: released
#11
DATE: 15/5/93
PRIORITY: 2
ISIS flexure test needs to be carried out. See note on how to do
this by Dave Carter. Data can be forwarded to Dave for reduction.
STATUS: 24/5/93: carried out okay on red and blue arm.
Horrible on red arm, okay on blue arm
RESULT: released, but needs fixing
#13
DATE: 15/5/93
PRIORITY: 2
ISIS blue arm needs ND filters.
STATUS: see #84
RESULT: dismissed
#14
DATE: 15/5/93
PRIORITY: 3
Are there enough arc lamps and spares ? Hans has asked to make an
invertory of lamps
STATUS: 2/6/93: Emilios invesitgated and supplied list.
RESULT: -
#15
DATE: 15/5/93
PRIORITY: 3
Get more channels for detectors going; currently 3 but 4 is minimum
STATUS: 26/1/95: awaiting hardware.
5/5/95: more available for some time now.
RESULT: done
#16
DATE: 15/5/93
PRIORITY: 1
writing high-time res. document for ISIS manual
STATUS: Vik working on it
RESULT: released
#17
DATE: 15/5/93
PRIORITY: 2
Have a look at ISIS manual in WHT control room.
STATUS: I don't like the fact that vital pol. info has disappeared from
the manual. Also, there is no pol manual in the control room anymore.
Have send Davce Carter relevant part of pol. manual to put into ISIS
manual.
RESULT: released
#18
DATE: 15/5/93
PRIORITY: 2
check fits headers are okay with new DAS version, in particular
check polarization headers.
STATUS: 26/7/93: VSD checked okay.
RESULT: released
#19
DATE: 15/5/93
PRIORITY: 2
get focus offset on AUX port sorted out so that capstans can be
set to the correct value, and so we know what the offset will be for
different filters which could then be coded in ICL to give automatic offset
STATUS: 5/5/95: done to certain extend. AUX box is being re-designed
so this will happen at some point.
RESULT: dismiss
#20
DATE: 15/5/93
PRIORITY: 2
ICL routine to be written and included in automatic startup
Chris Benn's routines (are already there),
focus routine James Scoby working on this (Hartmann test),
routine for polarization (4 frames with different settings).
May store in DMS data cube to reduce overheads, like in Steve Unger's
ICL procedure used with TAURUS and CCD. However, this may be more
difficult to handle in reduction packages.
STATUS: 25/7/93: VSD made routine for circular and linear pol obs.
16/11/93: Peter Fishwick working on focus routine.
11/3/94: Pete's FINEFOCUS routine appears to be working fine.
Reliability and agreement with hand-measured focus is checked
and found okay. FINEFOCUS is slow, but probably more accurate.
Also some of the error messages, questions and output are somewhat
confusing. The routine seems to handle hardware errors fine.
11/3/94: also we have now a routine to automatically carry out
a flexure test. This routine called FLEX still needs some tweaking
to make it more robust.
RESULT: released
#21
DATE: 15/5/93
PRIORITY: 2
make clear to the users when IPCS is the preferred detector on ISIS
STATUS: 25/7/93: will be put in Gemini
RESULT: released
#22
DATE: 15/5/93
PRIORITY: 2
what are the offsets in collimator position introduced by the
below-slit filters ?
STATUS: 19/12/93: in red arm the GG495 introduces a shift of
-0.40 pixels, and the 5400 dichroic gives and additional shift
of -0.20 pixels in the EEV detector (focus offsets are +800 and
+400 respectively).
RESULT: released
#23
DATE: 15/5/93
PRIORITY: 2
we need a GG395 filter for ISIS red arm. Chris Benn may have ordered
this; check. Also, find RFILTA slide; seems to have disappreared.
Why not get a couple of different GG filters.
Also for the blue arm we need ordersorting filters.
STATUS: 25/1/94: see item #84
RESULT: released
#24
DATE: 15/5/93
PRIORITY: 3
timing on data taken with ISIS needs to be improved drastically.
John McLean is working on this.
STATUS: 28/2/94: all done now. CCD controllers update automatically
once every day from time service.
RESULT: released
#25
DATE: 15/5/93
PRIORITY: 2
light leaks in ISIS. Chris Benn has ordered cloth, which needs
fitting when it arrives.
STATUS: now new shutter units are fitted with much imporved
light locks. Solves the problem for the CCD cameras.
(ISIS itself may need some improvement in reducing light leaks
but it is not really known how bad it is).
RESULT: released
#26
DATE: 15/5/93
PRIORITY: 2
TEK chip goes non-linear when used in binned mode.
can this be improved ?
STATUS: 22/8/93: binning spatially is okay; not in wavelenght (X)
28/2/94: this is all okay now.
RESULT: released
#27
DATE: 15/5/93
PRIORITY: 2
Harris filters need propper storage boxes; check with Hans Slingerland
STATUS: 15/5/93: Hans says he and Dave King are looking for then - monitor
1/12/95: Mariet looking into the general issue of storing filters.
RESULT: completed - all filters propperly stored
#28
DATE: 15/5/93
PRIORITY: 2
check out error in FORTH program on DMS to measure gain.
STATUS: 10/4/94: Daniel and I have compared the way we measure
the gain. These agree. Gain measurements can easily be done in
IRAF, so there is no real need for the DMS.
RESULT: released
#29
DATE: 15/5/93
PRIORITY: 2
measure fringing with TEK1 on ISIS red arm. Check with Paul Jorden
on how to do this.
STATUS: 17/7/93: images taken - fringing bad beyond ~6000 A;
ripples are about 2-3 percent.
11/8/93: eve further in red 9000-10000 fringing going up from
10 % to evenn 20 %.
RESULT: released
#30
DATE: 15/5/93
PRIORITY: 2
put TV focus and telescope focus offsets for polarization
optics in manual. Vik has measured these.
STATUS: 7/93: measured and documented
RESULT: released
#31
DATE: 15/5/93
PRIORITY: 2
add to ISIS status board the row number corresponding to the slit position.
STATUS: 26/1/95: status board will virtually disappear and all info
will appear on mimic screens. A narrow dekker image will immediate
tell what how the slit centre projects onto the CCD.
RESULT: dismissed
#32
DATE: 15/5/93
PRIORITY: 3
TV filter posn EMPTY is NOT ! empty but opaque
STATUS: 26/7/93: now it's possible to edit any of the optics using
the CHANGE command
RESULT: released
#33
DATE: 15/5/93
PRIORITY: 1
high time resolution observations can be done by storing data cubes
on DMS. This should be offered at some point as a standard option,
but needs software effort to do.
STATUS: ??/7/94: the TAURUS routines have been hacked and made into
a crude ISIS cube routine. This was tested, but not very extensively.
It was found that the overheads between exposures was still substantial
and not much less than when doing a multrun. Also header info is lost
for most exposured. It still is useful to try to get an optimized
CUBE routine.
5/5/95: Frank Gribbin and Daniel Mathews are looking into speeding
up our CCDs.
1/1/96: A fast multrun option has been identified, and on to of that
a project has been initiated to get really high-speed performance out of
our CCDs.
RESULT: 97: high-speed CCD drift mode implemented
#34
DATE: 15/5/93
PRIORITY: 2
check out data reduction software for FOS; check with Guy Rixon
STATUS: 12/4/94: a new IRAF reduction package has been written and
delivered. This is for quick-look results. FOS reduction can be
done with standard tools like IRAF.
RESULT: dismissed
#35
DATE: 24/5/93
PRIORITY: 2
positioning of grating on blue arm is off by +130 units
STATUS: 27/5/93: under investigation; Sue Worswick will look at this
6/6/93: given Dave King more details about this and they'll look
into the problem further. Dave says that the blue collimator is known
to be somewhat unreliably clamped.
8/7/93: look okay now: with 158 grating only few pixels off
RESULT: released
#36
DATE: 28/5/93
PRIORITY: 2
ripple in polarization spectrum reported using 1200 grating.
See Jaap Tinbergen's e-mail messages
STATUS: 28/5/93: got very good data on pol. standard star. This should
reveal it, if present. The pol. standard star indeed shows ripple with
amplitude dependent on wavelength, maximum of order 0.3 % !
Both amplitude and phase stable over different sets of observation.
Checked on zero pol. star and lamp through polarizer, same results.
Also R600R grating shows effect. Also clear intensity variations which
are probably related to pol. ripple.
Also did many checks on tungsten lamp. Jaap Tinbergen is investigating
the features. See ripple report for further details.
'gain' spectrum needs to be looked at.
1/11/94: we received a report (plots) that observers experienced very
similar efects at the AAT. It was also reported that the ripple can be
removed quite well by fitting a sine curve with variable frequency and
amplitude as a function of wavelength and correcting the data using
this fit. The resulting accuracy of the polarization measurement was
much better than the amplitude of the ripple.
RESULT: understood (more or less).
#37
DATE: 15/5/93
PRIORITY: 2
make dekker for pol. measurements which has bigger holes.
Request from Italian observers.
STATUS: 7/93: measurements made. Now need to specify in detail
for `final' dekker slide. Not urgent for Italian group
16/11/93: will be part of the new dekkers which will be ordered.
see job #58
RESULT: dismissed
#38
DATE: 15/5/93
PRIORITY: 2
think about usefulness to have red cutoff filter in ISIS (and UES)
see Jaap's e-mail about this
STATUS: 26/1/95: yes, this is still very useful indeed and hopefully
we can place an order soon. Faxed Schott to ask for more info on
their latest filters. It turns out that flat fielding quality in
the blue is most likely limited by red photons scattering into the
blue arm. A blue pass filter may cut this down. See # 84
RESULT: dismissed
#39
DATE: 15/5/93
PRIORITY: 2
determine position angle of optical axis of:
-1- calcite
-2- hw plate
-3- polarisation filters
STATUS: done: calcite axis is in same plane as slit since the throw of
the offset beam is in the spatial direction.
HW plate axis (at wavelength of 6960 A) was determined to be ~4 degrees
relative to the slit, measured from 2 pol. standard stars.
Polarizer #1 was calibrated using the fact that the HW plate has a 4
degrees offset (i.e. 8 degrees throw). Polarizer axis was estimated to
be 135 degrees. This was checked by eye and turns out to be correct.
RESULT: released
#40
DATE: 4/6/93
PRIORITY: 1
Flexure tests show very bad results for red arm; needs investigating.
Possibly the red fold is loose.
STATUS: turned out to be the due to capstan being bad. fixed.
Later flexure test shows better results, but still the red arm
has flexure which is about twice as large as the blue arm.
RESULT: released, but monitor
#41
DATE: 1/7/93
PRIORITY: 1
reported spatial shift of images taken with the TEK chip.
STATUS: first test to confirm this on 50 read-outs gave no shifts,
but second, similar test did show 11 images out of 200 shifted by 1 pixel.
EEV CCD is okay.
11/9/93: finally found a fauly board; replaced and problem gone.
RESULT: fixed
#42
DATE: 18/7/93
PRIORITY: 2
HW plate sticking and moving in beam when telescope is at
high airmass. Aparently the mimic does not give a warning that
it is not in place.
STATUS: 26/10/93: the problem was that the force acting on the HW
slide was difficult to control. Working against gravity it was too
weak and with gravity too strong. The piston has been changed so there
is a natural braking action and the strenght of the piston action will
always be large enough.
RESULT: released
#45
DATE: 19/7/93
PRIORITY: 1
TEK chip has high dark current (factor ~5 too high !) of about
100 e- /hour.
STATUS: 15/8/93: CCD has a long thermal response and must be kept
cold at all time. If not, dark current goes up as it is very sensitive
to temperature.
RESULT: understood
#46
DATE: 29/7/93
PRIORITY: 2
clean HW and QW plate as soon as the opportunity is there
STATUS: 25/10/93: done
RESULT: released
#43
DATE: 19/7/93
PRIORITY: 2
request from Paul Jorden to compare efficiency of ISEC TV in
CASS A&G with TEK chip in AUX f/11 beam. For further details
see note from PRJ
STATUS: 5/5/95: no request since - abandon
RESULT: dismiss
#48
DATE: 30/8/93
PRIORITY: 2
CCD read-out must be faster. Observers complaining and saying that
at the AAT it's going much faster than it is with us.
STATUS: 28/2/94: different readout speeds now available.
RESULT: released
#49
DATE: 1/8/93
PRIORITY: 2
RMS of calibrate in slit view is systematically better that in
direct view, suggesting flexure in direct view mirror
STATUS: 25/10/93: improved stiffness in mirror drive, but see
job #63 for more details
RESULT: released
#50
DATE: 1/10/93
PRIORITY: 2
autoguider travel must be brought back to accomodate additional
optics in the system
STATUS: 24/10/93: autoguider CCD housing has been brought forward
by about 2 mm. Optimal focus position in normal slit view is now 4400
with the clear autoguider filter.
It's scale is measured to be 4200 units per millimeter shift in
telescope focus (same sign).
RESULT: released
#53
DATE: 8/11/93
PRIORITY: 2
Request for Further ISIS Arc Maps. ---> see job #9
STATUS: 29/6/94: identical to job #9
RESULT: released
#54
DATE: 6/11/93
PRIORITY: 2
when inspecting the filter movement of the below slit filters in
the blue arm it was seen, looking through the grating port, that
when a long filter slide is used in the blue arm vignetting may
occur.
STATUS: 8/11/93: checked vignetting caused by the long filter slide
in the blue arm and it was found to be only of the order of 1-2 %
with the slide out of the beam.
11/11/93: Sue showed me that the main distinction between the filters
in the blue arm and the ones in the red arm is that the blue ones
are closer to the slit and hence are more narrow than the red ones.
RESULT: understood
#55
DATE: 9/11/93
PRIORITY: 2
measure response of o and e beam for different angles
of the HW and QW plates.
STATUS: 19/12/93: done a number of times ! This turns out to be
a reliable method to determine the action and angle positioning
accuracy of the wave plates.
RESULT: released
#58
DATE: 16/11/93
PRIORITY: 2
old job: write up specification for new set of dekker slides
for polarimetry.
STATUS: 16/11/93: measurements have been collected and we know
what we want.
19/12/93: specification written and send off to Cambridge for
further action and manufacturing. Dekker are manufacture in Roden.
RESULT: released
#59
DATE: 16/11/93
PRIORITY: 2
check out optics for imaging polarimetry; test setup; sort
out below slit filter slides and measure propper filter sizes.
Test the system thoroughly on the sky so we can support it.
STATUS: 16/11/93: done some trial exposures on lamps.
Charles Jenkins has been given the sizes and he's going to
order special filters for imaging polarimetry.
28/2/94: more test have been done, offsets determined and
acquisition worked out. Details will be documented.
RESULT: released
#61
DATE: 28/12/93
PRIORITY: 1
camera mirrors:
* are very dirty
* are loose, specially the blue one. The blue arm was showing
shifts of ~8 pixels with a flexure test. This is most likely
due to this mirror being loose. Also the lens in front of
the CCD on the blue arm feels loose, although it is spring
loaded held against its mounting.
STATUS: tightened the camera mirrors. Problem is that the screws
cannot be secured and might work loose over time.
Cleaned one mirror in the red arm and one mirror in the blue arm.
Mirror in red arm looks okay, apart from mark in lower right corner.
Mirror in blue arm left streaks after cleaning with wet cloth.
Edges of this mirror look poor.
Also cleaned one (accessible) surface of each camera lens - fine.
A new flexure test still showed too large a shift in the blue arm.
CCD lenses were cleaned and tightened.
talk with Dave Jackson or Sue about giving them a professional clean.
5/5/95: cleaning and re-alignment work is happening now.
RESULT: COMPLETED
#64
DATE: 10/3/94
PRIORITY: 2
reported poor spatial image with IPCS.
STATUS: 10/3/94: it turned out that the collimator setting was
11000 while the optimum value is about 6000. The spatial width
of the narrow dekker is ~4.5 pixels on the TEK CCD at coll posn
6500, while at 11000 it is ~7.5 pixels.
We need to establish how the spatial profile degrades with
collimator position.
11/3/94: the following FWHM values were measured for the spatial profile
in the red and blue arm for different collimator positions with the
tungsten lamp illuminating the 1.2" narrow dekker (in units of pixels with
EEV on the red arm and TEK on the blue arm. Cenwave in red was 6500 A
and in blue 6000 A):
coll posn. red blue
---------------------------------
0 4.5
1000 8.2 4.0
2000 6.7 3.9
3000 5.4 3.8
4000 4.7 3.9
5000 4.2 3.9
6000 4.1 4.1
7000 4.1 4.5
8000 3.9 5.3
9000 3.9 6.2
10000 4.1 6.9
11000 4.4 7.7
12000 5.6 8.6
13000 6.5
14000 8.3
Hence, the optimal value for the red arm is 8000 with an allowed
range from 3500 to 11500, and in the blue arm the optimum is 4000
with an allowed range of 0 to 7500 units.
But, note that this is not really a good test since the range
depends on the actual position of the detector. A propper test
would involve changing the detector position, refocussing by changing
the collimator position, and then measuring the width of the
special profile.
RESULT: released
#65
DATE: 11/3/94
PRIORITY: 1
new dichroics are delivered and mounted
Dave King writes:
I have set up all the new dichroics for ISIS now. All the old ones on
thin glass have been retired. We now have:--
Slide pos 0 pos 1 pos 2
1 clear flat 7500
2 clear flat 6100
3 clear flat 5700
4 clear flat 5400
5 clear prism 4500
I have also left spares of the 5700,5400 and 4500 dichroics, and a
couple of spare aluminised flats.
The mountings fixing the castings to the base plates are affected greatly
by the tension on the fixing bolts. It should not be necessary to remove
the castings in normal use. If for some reason you do want to take the
glassware out, in most cases it is possible to do so by just removing the
sprung metal clips holding the glass in the castings. If not, them remove the
fixing bolts carefully, noting the tightness to which they are turned
beforehand. I think it willbe almost impossible to remove and replace the
casting from the base plates and maintain the alignment.
Transmission and reflectivity needs to be measured.
STATUS: 1/4/94
done for 4 of the dichroics. Transmission is easy to measure;
reflectivity is more tricky as it is done relative to the reflectivity
of a mirror and in the blue the tungsten lamp is faint and stray light
ruines the measurement.
Re-done the transmission measurements since the tungsten lamp measurements
were not good due to scattered light. Used blue star instead and this time
it worked !
RESULT: released
#66
DATE: 22/3/94
PRIORITY: 2
check how CCD pixels are sensitive to the o or e beam for
flat fielding purposes
STATUS: 22/3/94: took many flat fields with polarising filters in
the beam. The 2 pol. filters give the correct angle of polarisation
to fit with the o and e ray of the calcite plate.
The flat fields were taken only with the pol. in the beam (no calcite
or waveplate). The summed flat field for each yielded more than 10^6
photons per pixel.
The RMS on each summed flat field was of the order of 3 %, as expected
for normal pixel-to-pixel variations.
By dividing flat fields with filter by the flat field without the filter
(i.e. by flat-fielding the pol. filter flat field) the RMS went down to
0.25 %. The RMS on the ratio of the flat fields with both pol filters
was 0.35 %. This was all done with the EEV CCD.
Hence, in polarisation mode (i.e. with the calcite) the pixel-to-pixel
variations are gouverned by normal flat field effect. Only at a level
~0.3 % polarisation effect are becoming important.
RESULT: released
#67
DATE: 10/4/94
PRIORITY: 2
Measure the dependence of the width of the spatial profile as a function
of collimator position. This must be done by setting the capstan screws
of the detector to a series of different values, focussing the spectrograph
at each position by shifting the collimator, and measuring the width of
the spatial profile of the narrow dekker. See for preliminary results #64
STATUS: 6/5/95: work on servicing and re-alignment of ISIS optics
should give us new optimal autocollimation values.
The optimal values reported by Sue W are:
5687 in the blue arm and 9603 in the red arm.
RESULT: released
#69
DATE: 1/5/94
PRIORITY: 1
the ISIS shutters have always been unreliable. A new design has been
made and tested, based on a comercially available shutter. The prototype
has been operated for many days continuously without problems.
New shutter plates are being made so that both arms will have the new
shutters fitted, and spares will be available for quick emergency
changes in case of failure.
STATUS: 10/7/94: new shutters have been fitted and have been operating
without failure for months now. More shutters will be made
RESULT: released
#70
DATE: 10/4/94
PRIORITY: 2
some more work has been done to improve the performance of the HWP
mechanism. The gearing has been machined more accurately, an extra
zero-point sensor was wired up as backup (identical to the working
sensor), an additional (different) sensor was tested but not fitted
because of space problems (this may improve the zero set accuracy),
and the mechanical rotational arrangement from the QWP was swapped
with that of the HWP (QWP movement was better than HWP movement; HWP
is mainly used).
STATUS: tested now, but monitor
RESULT: released
#74
DATE: 2/6/94
PRIORITY: 2
one of the TV cables (used for the fibres TV) in the CASS cable
wrap is broken.
STATUS: -
RESULT: fixed
#75
DATE: 4/6/94
PRIORITY: 2
dekker appears to move on TV - flexure ?
STATUS: Yes, must be the same flexure which is causing the poor
calibrate results. This is under investigation as part of item 63.
RESULT: in-hand
#76
DATE: 5/6/94
PRIORITY: 2
is NEWTIME command working properly ?
STATUS: 7/6/94: tested: answer is "NO", only under certain circumstances
it is working. Report presented to Pete Fishwick:
* When the NEWTIME command is used to set the exposure time
of a current exposure to a value which is shorter than the
orriginally set time, but still longer that the actually
exposed time [ T(actual) < T(new) < T(set) ] the new time is
handeled correctly; the exposure stops at T(new) and the log
and FITS headers carry the correct time.
However, the shutter closes and opens after issuing the command,
and I've seen that the shutter closes and stays closed, or
remains open from the start, as it should. Clearly the NEWTIME
command should not affect the shutter in this case.
Furthermore, the `set time' on the mimic does not update to the
new value, so the user is not notified that the command was
accepted.
* When the exposure time is set to a value which is shorter than
the actually exposed time [ T(new) < T(actual) < T(set) ]
the exposure stops immediately, the CCD is read out, and the log
and FITS headers carry the correct actually exposed time.
* When the exposure time is set to a value which is longer than
the set time [ T(actual) < T(set) < T(new) ], again the shutter
closes for an instant after issuing the command, but not each
time I've tried it. NEWTIME should not affect the shutter in
this case.
The strangest thing is this mode is that when the NEWTIME command
is issued in the first ~30 sec of the run it carries on for
some 25 sec. and then suddenly the exposed time jumps to the
new time and the CCD read out !! In this case the observer thinks
the exposure has finished normally with the newly set time while
it actually has not; the log and FITS headers all lie.
Furthermore, the `set time' on the mimic does not update to
the new value, so the user is not notified that the command
was accepted.
These features are identical on the red and blue arms of ISIS.
We must not used NEWTIME until these problems are fixed.
7/11/94: in newest software release (version 9.0) this is fixed
RESULT: released
#77
DATE: 1/6/94
PRIORITY: 2
test ISIS focus at different positions of the telescope
STATUS: 27/6/94: tested for mount PA -90, 0, 90 180 and
telescope at EL 90, 60 and 30; measured hartmann shift
at each position, in red and blue arm. Maximum (differential)
shift was 0.33 in red and 0.38 pixels in blue (EEV6 on red
arm and TEK1 on blue arm). Sigma in shifts: 0.11 pixels in
both red and blue arm.
RESULT: released
#78
DATE: 1/6/94
PRIORITY: 2
we need to know the response curve of the R1200B grating
STATUS: 28/6/94: asked Dave King
7/11/94: found response curve documented in ISIS Astronomers guide.
Also updated ISIS manual in control room.
RESULT: released
#80
DATE: 25/7/94
PRIORITY: 2
new ISIS shutters are now fitted on the red and blue arm.
These not only contain a comercial shutter (piston driven by
gaseous nitrogen), but also a much improved light baffle which
is spring loaded against the CCD front flange.
Tests need to be carried out to check for residual light leaks
and the shutter performance needs to be measured for different
levels of the nitrogen pressure.
STATUS: 3/8/94
light baffle: superb ! with lights on in the dome and even with
tungsten lamp on in ISIS, no light detected; just dark current
on 1000 sec. exposure in the blue as well as red arm.
Shutter reliability from multiple exposures was found to be better
that 0.1 % for a 1 sec. exposure time.
Photometric accuracy over full field (i.e. including shutter timing
delays etc.) is 1% for a 1 sec. exposure.
Vignetting by shutter wings: 0.2% from center to corner for 1 sec.
exposure.
More details are documented elsewhere.
Nitrogen pressure was set at 3 bar (standard)
RESULT: released
#83
DATE: 25/1/95
PRIORITY: 1
IRAF based ISIS setup scripts need to be written for
tilt, rotation and focussing of the spectrograph.
STATUS: the routines have been written and tested okay.
They will be improved when more experiece developes.
RESULT: completed
#85
DATE: 16/12/93
PRIORITY: 3
Contribution to the ISIS change manual from the astronomy side.
STATUS: 26/1/95: completed
RESULT: completed
#86
DATE: 22/2/95
PRIORITY: 1
investigate possibility of installing an optical CCD in
the Hitchiker port.
STATUS: 5/5/95: not the optimal option. Better would be to
have keep the optical CCD where AUX port is, and an IR array
opposite to AUX. One mirror then has be be flipped. See #87.
The Hitchhker port needs re-imaging optics and is quit far
off axis which degrades the image quality.
RESULT: abandon
#92
DATE: 6/5/95
PRIORITY: 1
following the ISIS servicing and re-alignment (May 95) we must
re-measure the dichroic transmission with great accuracy and
covering the full wavelength range.
STATUS: 1/12/95
RESULT: done
#93
DATE: 6/5/95
PRIORITY: 1
following the ISIS servicing and re-alignment (May 95) we must
re-establish the system efficiency.
STATUS: 1/12/95
RESULT: done
#94
DATE: 6/5/95
PRIORITY: 1
investigate light tightness of ISIS structure.
STATUS: 6/5/95: Jaap Haan carefully looked at light leaks and
found several spots where light could get in. These have all
been repaired.
RESULT: released
#98
DATE: 10/7/97
PRIORITY: 1
determine focus offsets for dichroics
STATUS: 10/7/97
completed. Focus offsets for red arm are:
dichroic focus offset
4500 1350
5400 1390
5700 1350
6100 1430
7500 1340
RESULT: completed
#101
DATE: 8/4/98
PRIORITY: 1
Check the Operational characteristics of the EEV10a chip when binned,
and also in different read-out modes. Especially check the perfomance
when binned both spatially and spectrally as this will be an attractive
(and perhaps default!) mode for observers to use.
Check also flux-standards for the EEV10a.
STATUS: Done
RESULT: The EEV10a chip has been setup incorrectly. Moved to IDS, and
Simon Tulloch will investigate in April/May 1999.
#102
DATE: 8/4/98
PRIORITY: 2
Perform a new ISIS focus-flexure test, and use the data to re-measure the
normal flexure in the instrument. Peter Sorenson would like to help
out with this as part of his daytime duty requirements
STATUS: Done SJS and PMS performed test on 7/4/98, results available.
RESULT:
#104
DATE: 15/4/98
PRIORITY: 1
Complete ISIS web-pages to an Observer acceptable format and link to
the ING/WHT main pages.
STATUS: Done, but ungoing
RESULT:
#105
DATE: 1/8/98
PRIORITY: 1
Commissioning of new EEV12 chip for ISIS blue
STATUS: Done, and all characteristics released
RESULT: Excellent performance, up to spec and expectations
Stephen Smartt (ISIS Instrument Specialist) sjst@ing.iac.es